# Probing anomalous Higgs boson couplings in Higgs plus jet production at NLO QCD with full $m_{t}$-dependence

### v1

@@ 3 Numerical results and discussion of anomalous couplings
The results presented in this section were obtained using the PDF4LHC15_nlo_30_pdfas parton distribution functions Butterworth:2015oua ; Dulat:2015mca ; Harland-Lang:2014zoa ; NNPDF:2014otw interfaced to our code via LHAPDF Buckley:2014ana , along with the corresponding value for $\alpha_{s}$. The masses of the Higgs boson and the top quark have been fixed, as in the virtual amplitude, to $m_{H}=125~{}\mathrm{G}\mathrm{eV}$, $m_{t}=173.055~{}\mathrm{G}\mathrm{eV}$, respectively. Their widths have been set to zero. Jets are clustered with the anti-$k_{T}$ algorithm Cacciari:2008gp as implemented in the fastjet package Cacciari:2005hq ; Cacciari:2011ma , with jet radius $R=0.4$ and a minimum transverse momentum $p_{T,\rm{min}}^{\rm{jet}}=30~{}\mathrm{G}\mathrm{eV}$. The central scale is given by
@@ 3.1 Total cross sections and heat maps
The total cross sections obtained with the above settings and for different values of the minimum transverse momentum of the Higgs boson, $p_{T,H}^{\rm{min}}$, are given in Table 1, where we compare the benchmark point $c_{t}=0.9,c_{g}=1/15$ to the SM case. We can clearly see that at LO, the difference between BSM and SM is within the corresponding scale uncertainties. At NLO, the scale uncertainties are reduced and the difference becomes noticeable for highly boosted Higgs bosons, but only for a $p_{T,H}$ cut of $800\,\mathrm{G}\mathrm{eV}$ the difference is clearly outside the scale uncertainties for the considered benchmark point. This shows that already for small deviations from the SM there can be a measurable difference, if the Higgs boson is very highly boosted. Going away further from the SM values for $c_{g}$ and $c_{t}$, the difference would become more pronounced and start at a smaller $p_{T,H}$–cut values. This will become apparent in the discussion of the heat maps shown in Figs. 4 and 5. In Table 2 we show the ratio of the HEFT at the benchmark point $c_{t}=0.9$, $c_{g}=1/15$ to the SM case, including Monte Carlo uncertainties rather than scale uncertainties, as the scale uncertainties are correlated between HEFT and the SM and therefore largely cancel. The ratios clearly demonstrate the behaviour discussed above, i.e. the difference between HEFT and the SM becoming more pronounced for higher values of $p_{T,H}^{\rm{min}}$.
@@
0 $(9\begin{subarray}{c}+3\\ -2\end{subarray})\cdot 10^{3}$ $(15\begin{subarray}{c}+0\\ -2\end{subarray})\cdot 10^{3}$ $(9\begin{subarray}{c}+3\\ -2\end{subarray})\cdot 10^{3}$ $(15\begin{subarray}{c}+0\\ -2\end{subarray})\cdot 10^{3}$
50 $(4.6\begin{subarray}{c}+0.2\\ -0.1\end{subarray})\cdot 10^{3}$ $(9.3\begin{subarray}{c}+0.2\\ -1.7\end{subarray})\cdot 10^{3}$ $(4.6\begin{subarray}{c}+0.2\\ -0.1\end{subarray})\cdot 10^{3}$ $(9.3\begin{subarray}{c}+0.0\\ -1.7\end{subarray})\cdot 10^{3}$
100 $(1.4\begin{subarray}{c}+0.6\\ -0.4\end{subarray})\cdot 10^{3}$ $(2.84\begin{subarray}{c}+0.04\\ -0.5\end{subarray})\cdot 10^{3}$ $(1.4\begin{subarray}{c}+0.6\\ -0.3\end{subarray})\cdot 10^{3}$ $(2.88\begin{subarray}{c}+0\\ -0.5\end{subarray})\cdot 10^{3}$
200 $(2.2\begin{subarray}{c}+1.0\\ -0.6\end{subarray})\cdot 10^{2}$ $(4.6\begin{subarray}{c}+0.2\\ -0.9\end{subarray})\cdot 10^{2}$ $(2.2\begin{subarray}{c}+0.9\\ -0.6\end{subarray})\cdot 10^{2}$ $(4.6\begin{subarray}{c}+1\\ -9\end{subarray})\cdot 10^{2}$
400 $13\begin{subarray}{c}+6\\ -4\end{subarray}$ $28\begin{subarray}{c}+2\\ -6\end{subarray}$ $12\begin{subarray}{c}+5\\ -3\end{subarray}$ $25\begin{subarray}{c}+1.5\\ -5\end{subarray}$
600 $1.6\begin{subarray}{c}0.7\\ -0.5\end{subarray}$ $3.3\begin{subarray}{c}+0.4\\ -0.7\end{subarray}$ $1.3\begin{subarray}{c}+0.6\\ -0.4\end{subarray}$ $2.7\begin{subarray}{c}+0.2\\ -0.5\end{subarray}$
800 $0.28\begin{subarray}{c}+0.13\\ -0.08\end{subarray}$ $0.57\begin{subarray}{c}+0.07\\ -0.11\end{subarray}$ $0.20\begin{subarray}{c}+0.10\\ -0.06\end{subarray}$ $0.41\begin{subarray}{c}+0.04\\ -0.08\end{subarray}$
Table 1: Total cross sections for different values of $p_{T,H}^{\rm{min}}$, The main value is based on the central scale and the uncertainties are obtained from three-point variations around the central scale. The HEFT values are for the benchmark point $c_{t}=0.9,c_{g}=1/15$.
@@
where the last relation is based on the fact that, if we denote the effective gluon-Higgs coupling in the heavy-top-limit of the SM as $c_{g,\mathrm{HTL}}$, then the cross section can be written as $\sigma_{\rm{HTL}}=c_{g,\rm{HTL}}^{2}C_{\rm{HTL}}$, with $c_{g,\rm{HTL}}=2/3$ as explained above. Thus it is sufficient to compute the cross section for different values of $c_{g}$ and $c_{t}$ and then fit the coefficients $A,B$ and $C$. For the fit we chose some ($c_{t},c_{g}$) value pairs across a wide range, even outside the experimental limits, in order to guarantee a good fit.
@@
In Table 3 we list the fitted values for the coefficients using different cuts on the Higgs boson transverse momentum. The corresponding values for variations around the central scale are given in appendix A.
We see that for all cuts the largest coefficient is always the $C$-coefficient, followed by $B$ and then $A$. Thus the dominant contributions to the cross sections stem from purely HTL-like diagrams. Note that the reason why for the chosen benchmark point the SM-like diagrams are more important comes from the fact that $c_{t}=0.9\gg c_{g}=1/15$, compatible with current constraints. Furthermore, the higher the cut on the transverse momentum of the Higgs boson, the bigger the ratios $C/A$ and $B/A$ become, indicating again that the deviations from the SM start to become more important for high $p_{T,H}$. This stems from the fact that $d\sigma/dp_{T,H}\sim 1/p_{T,H}^{a}$ scales with $a=2$ in the full theory and with $a=1$ in the HTL Caola:2016upw . We can also see that $C_{\mathrm{HTL}}\approx C$ for all values of the cut. This is to be expected since the heavy top limit in the SM corresponds to the HEFT with $c_{g}=2/3$, excluding those top-loops where the top quark only couples to gluons. The latter are suppressed in the heavy top limit since they do not involve a Yukawa coupling. Thus this gives a cross-check of our computations.
@@
0 $15.4\pm 0.2$ $21.3\pm 1.6$ $30.6\pm 2.8$ $30.7\pm 0.2$
50 $9.35\pm 0.13$ $13.4\pm 0.3$ $19.2\pm 0.6$ $18.9\pm 0.4$
100 $2.88\pm 0.04$ $4.16\pm 0.06$ $6.15\pm 0.1$ $6.10\pm 0.09$
200 $(4.57\pm 0.03)\cdot 10^{-1}$ $(7.22\pm 0.09)\cdot 10^{-1}$ $(11.99\pm 0.15)\cdot 10^{-1}$ $(12.1\pm 0.3)\cdot 10^{-1}$
400 $(2.54\pm 0.01)\cdot 10^{-2}$ $(5.61\pm 0.08)\cdot 10^{-2}$ $(13.6\pm 0.1)\cdot 10^{-2}$ $(13.6\pm 0.4)\cdot 10^{-2}$
600 $(2.66\pm 0.01)\cdot 10^{-3}$ $(8.19\pm 0.11)\cdot 10^{-3}$ $(28.3\pm 0.2)\cdot 10^{-3}$ $(28.2\pm 0.3)\cdot 10^{-3}$
800 $(4.14\pm 0.02)\cdot 10^{-4}$ $(16.8\pm 0.2)\cdot 10^{-4}$ $(78.3\pm 0.5)\cdot 10^{-4}$ $(77.7\pm 0.2)\cdot 10^{-4}$
@@ 3.2 Higgs boson transverse momentum distributions
The $p_{T,H}$ distribution with a minimum $p_{T}$-cut of $30\,\mathrm{G}\mathrm{eV}$ on the jet is shown in Fig. 6, for the SM, the HTL and two HEFT benchmark points. Up to values of $p_{T,H}\approx 300~{}\mathrm{G}\mathrm{eV}$ there is no significant difference between the considered predictions, but for larger $p_{T,H}$ values the heavy top limit shows large deviations from the SM, whereas the two HEFT parameter points lead to significantly smaller deviations. Thus it is very important to use the SM predictions with full $m_{t}$-dependence, otherwise the approximation given by the HTL could mask enhancements in the tail which are in fact BSM effects.
@@
Figure 6: Higgs boson transverse momentum distribution for two HEFT benchmark points, $(c_{t},c_{g})=(0.9,1/15)$ and $(c_{t},c_{g})=(1.1,-1/15)$, compared to the SM case and to the heavy top limit. The bands denote 3-point scale variations around the central scale $\mu_{0}=H_{T}/2$.
@@ 4 Conclusions
We have presented results for Higgs boson production in association with one jet, combining full NLO QCD corrections with the leading operators within Higgs Effective Field Theory (HEFT), which can modify the top-Higgs Yukawa coupling and induce an effective Higgs-gluon coupling. We have taken into account constraints arising from the fact that inclusive Higgs production measurements show very good agreement with the SM prediction, assuming that subleading operators, such as the chromomagnetic operator or four-fermion operators, do not play a substantial role. We found that there are combinations of $c_{t}$ and $c_{g}$ that reproduce the inclusive Higgs production cross section at the 10% level, while changing the Higgs + jet cross section to an extent that exceeds the scale uncertainties in the highly boosted Higgs regime, i.e. for $p_{T,H}^{\mathrm{min}}\gtrsim 600~{}\mathrm{G}\mathrm{eV}$. We also found that the NLO K-factor can vary by more than 30% compared to the SM K-factor within the allowed range of $c_{t}$ and $c_{g}$. Furthermore, we showed that including the full top quark mass dependence is important to avoid that the approximation given by the heavy top limit, leading to an enhancement in the tail of the $p_{T,H}$ distribution, is masking a BSM effect.
It remains to be investigated in more detail whether top mass renormalisation scheme uncertainties could swamp the effects of anomalous couplings even for highly boosted Higgs bosons; the results presented in Ref. Bonciani:2022jmb for the SM case suggest that the ratio between on-shell and $\overline{\mathrm{MS}}$ renormalisation schemes for the top quark is rather flat for large $p_{T,H}$ values, while the BSM effects grow rapidly with $p_{T,H}$. Similar considerations hold for QCD corrections beyond NLO, as the K-factors are close to constant in the large $p_{T,H}$ range. This holds for the NLO K-factors both in the HTL and in the full SM, as well as for KNNLO in the HTL Becker:2020rjp . The effects of subleading operators such as the chromomagnetic or four-fermion operators also deserve further study, as well as electroweak corrections.
We plan to publish the Monte Carlo code underlying our results within the framework of the POWHEG-BOX-V2, after having constructed a grid and an interpolation framework for the virtual 2-loop amplitudes.
@@ A Appendix
0 $15.4\pm 0.2$ $12.8\pm 0.2$ $14.4\pm 0.4$
50 $9.35\pm 0.13$ $7.68\pm 0.09$ $9.37\pm 0.24$
100 $2.88\pm 0.04$ $2.36\pm 0.02$ $2.86\pm 0.07$
200 $(4.57\pm 0.03)\cdot 10^{-1}$ $(3.71\pm 0.02)\cdot 10^{-1}$ $(4.70\pm 0.05)\cdot 10^{-1}$
400 $(2.54\pm 0.01)\cdot 10^{-2}$ $(2.01\pm 0.01)\cdot 10^{-2}$ $(2.69\pm 0.02)\cdot 10^{-2}$
600 $(2.66\pm 0.01)\cdot 10^{-3}$ $(2.12\pm 0.01)\cdot 10^{-3}$ $(2.87\pm 0.02)\cdot 10^{-3}$
800 $(4.14\pm 0.02)\cdot 10^{-4}$ $(3.31\pm 0.02)\cdot 10^{-4}$ $(4.50\pm 0.03)\cdot 10^{-4}$
@@
0 $21.3\pm 1.6$ $18.3\pm 0.4$ $22.8\pm 1.0$
50 $13.4\pm 0.3$ $11.0\pm 0.2$ $14.8\pm 0.6$
100 $4.16\pm 0.06$ $3.40\pm 0.04$ $4.60\pm 0.11$
200 $(7.22\pm 0.09)\cdot 10^{-1}$ $(5.85\pm 0.06)\cdot 10^{-1}$ $(8.10\pm 0.15)\cdot 10^{-1}$
400 $(5.61\pm 0.08)\cdot 10^{-2}$ $(4.50\pm 0.05)\cdot 10^{-2}$ $(6.43\pm 0.13)\cdot 10^{-2}$
600 $(8.19\pm 0.11)\cdot 10^{-3}$ $(6.56\pm 0.08)\cdot 10^{-3}$ $(9.47\pm 0.18)\cdot 10^{-3}$
800 $(16.8\pm 0.2)\cdot 10^{-4}$ $(13.4\pm 0.2)\cdot 10^{-4}$ $(19.5\pm 0.3)\cdot 10^{-4}$
@@
0 $30.6\pm 2.8$ $26.4\pm 0.6$ $37.3\pm 1.6$
50 $19.2\pm 0.6$ $15.8\pm 0.4$ $23.4\pm 1.0$
100 $6.15\pm 0.1$ $5.01\pm 0.07$ $7.48\pm 0.18$
200 $(11.99\pm 0.15)\cdot 10^{-1}$ $(9.76\pm 0.10)\cdot 10^{-1}$ $(14.4\pm 0.3)\cdot 10^{-1}$
400 $(13.6\pm 0.1)\cdot 10^{-2}$ $(11.2\pm 0.1)\cdot 10^{-2}$ $(16.4\pm 0.2)\cdot 10^{-2}$
600 $(28.3\pm 0.2)\cdot 10^{-3}$ $(23.0\pm 0.1)\cdot 10^{-3}$ $(34.2\pm 0.3)\cdot 10^{-3}$
800 $(78.3\pm 0.5)\cdot 10^{-4}$ $(63.2\pm 0.3)\cdot 10^{-4}$ $(95.0\pm 0.7)\cdot 10^{-4}$

### v2

@@ 3 Numerical results and discussion of anomalous couplings
The results presented in this section were obtained using the PDF4LHC21_40_pdfas parton distribution functions Butterworth:2015oua ; Dulat:2015mca ; Harland-Lang:2014zoa ; NNPDF:2014otw interfaced to our code via LHAPDF Buckley:2014ana , along with the corresponding value for $\alpha_{s}$. The masses of the Higgs boson and the top quark have been fixed, as in the virtual amplitude, to $m_{H}=125~{}\mathrm{G}\mathrm{eV}$, $m_{t}=173.055~{}\mathrm{G}\mathrm{eV}$, respectively. Their widths have been set to zero. Jets are clustered with the anti-$k_{T}$ algorithm Cacciari:2008gp as implemented in the fastjet package Cacciari:2005hq ; Cacciari:2011ma , with jet radius $R=0.4$ and a minimum transverse momentum $p_{T,\rm{min}}^{\rm{jet}}=30~{}\mathrm{G}\mathrm{eV}$. The central scale is given by
@@ 3.1 Total cross sections and heat maps
The total cross sections obtained with the above settings and for different values of the minimum transverse momentum of the Higgs boson, $p_{T,H}^{\rm{min}}$, are given in Table 1, where we compare the benchmark point $c_{t}=0.9,c_{g}=1/15$ to the SM case. We can clearly see that at LO, the difference between BSM and SM is within the corresponding scale uncertainties. At NLO, the scale uncertainties are reduced and the difference becomes noticeable for highly boosted Higgs bosons, but only for a $p_{T,H}$ cut of $800\,\mathrm{G}\mathrm{eV}$ the difference is clearly outside the scale uncertainties for the considered benchmark point. This shows that already for small deviations from the SM there can be a measurable difference, if the Higgs boson is very highly boosted. Going away further from the SM values for $c_{g}$ and $c_{t}$, the difference would become more pronounced and start at a smaller $p_{T,H}$–cut values. This will become apparent in the discussion of the heat maps shown in Figs. 4 and 5. In Table 2, we show the ratio of the HEFT at the benchmark point $c_{t}=0.9$, $c_{g}=1/15$ to the SM case, including Monte Carlo uncertainties rather than scale uncertainties, as the scale uncertainties are correlated between HEFT and the SM and therefore largely cancel. The ratios clearly demonstrate the behaviour discussed above, i.e. the difference between HEFT and the SM becoming more pronounced for higher values of $p_{T,H}^{\rm{min}}$.
@@
0 $(8\begin{subarray}{c}+3\\ -2\end{subarray})\cdot 10^{3}$ $(15\begin{subarray}{c}+0\\ -2\end{subarray})\cdot 10^{3}$ $(8\begin{subarray}{c}+3\\ -2\end{subarray})\cdot 10^{3}$ $(15\begin{subarray}{c}+0\\ -2\end{subarray})\cdot 10^{3}$
50 $(4.5\begin{subarray}{c}+1.8\\ -1.2\end{subarray})\cdot 10^{3}$ $(9.1\begin{subarray}{c}+0.1\\ -1.6\end{subarray})\cdot 10^{3}$ $(4.5\begin{subarray}{c}+1.8\\ -1.2\end{subarray})\cdot 10^{3}$ $(9.1\begin{subarray}{c}+0.0\\ -1.4\end{subarray})\cdot 10^{3}$
100 $(1.4\begin{subarray}{c}+0.6\\ -0.4\end{subarray})\cdot 10^{3}$ $(2.85\begin{subarray}{c}+0.08\\ -0.5\end{subarray})\cdot 10^{3}$ $(1.4\begin{subarray}{c}+0.6\\ -0.4\end{subarray})\cdot 10^{3}$ $(2.85\begin{subarray}{c}+0.01\\ -0.5\end{subarray})\cdot 10^{3}$
200 $(2.2\begin{subarray}{c}+0.9\\ -0.6\end{subarray})\cdot 10^{2}$ $(4.6\begin{subarray}{c}+0.2\\ -0.9\end{subarray})\cdot 10^{2}$ $(2.2\begin{subarray}{c}+0.9\\ -0.6\end{subarray})\cdot 10^{2}$ $(4.5\begin{subarray}{c}+0.1\\ -0.9\end{subarray})\cdot 10^{2}$
400 $13\begin{subarray}{c}+6\\ -4\end{subarray}$ $28\begin{subarray}{c}+2\\ -6\end{subarray}$ $12\begin{subarray}{c}+5\\ -3\end{subarray}$ $25\begin{subarray}{c}+1.6\\ -5\end{subarray}$
600 $1.6\begin{subarray}{c}0.7\\ -0.5\end{subarray}$ $3.3\begin{subarray}{c}+0.3\\ -0.7\end{subarray}$ $1.3\begin{subarray}{c}+0.6\\ -0.4\end{subarray}$ $2.7\begin{subarray}{c}+0.2\\ -0.5\end{subarray}$
800 $0.29\begin{subarray}{c}+0.14\\ -0.09\end{subarray}$ $0.60\begin{subarray}{c}+0.07\\ -0.12\end{subarray}$ $0.21\begin{subarray}{c}+0.10\\ -0.06\end{subarray}$ $0.43\begin{subarray}{c}+0.04\\ -0.09\end{subarray}$
Table 1: Total cross sections for different values of $p_{T,H}^{\rm{min}}$, The main value is based on the central scale and the uncertainties are obtained from three-point variations around the central scale. The HEFT values are for the benchmark point $c_{t}=0.9,c_{g}=1/15$. All cross sections have a minimal jet-$p_{\mathrm{T}}$ of $30~{}\mathrm{G}\mathrm{eV}$ and thus at LO, the Higgs boson also has a minimal transverse momentum of $30~{}\mathrm{G}\mathrm{eV}$.
@@
where the last relation is based on the fact that, if we denote the effective gluon-Higgs coupling in the HTL of the SM as $c_{g,\mathrm{HTL}}$, then the cross section can be written as $\sigma_{\rm{HTL}}=c_{g,\rm{HTL}}^{2}C_{\rm{HTL}}$, with $c_{g,\rm{HTL}}=2/3$ as explained above. Thus it is sufficient to compute the cross section for different values of $c_{g}$ and $c_{t}$ and then fit the coefficients $A,B$ and $C$. For the fit we chose some ($c_{t},c_{g}$) value pairs across a wide range, even outside the experimental limits, in order to guarantee a good fit.
@@
In Table 3, we list the fitted values for the coefficients using different cuts on the Higgs boson transverse momentum. The corresponding values for variations around the central scale are given in appendix A.
We see that for all cuts the largest coefficient is always the $C$-coefficient, followed by $B$ and then $A$. Thus the dominant contributions to the cross sections stem from purely HTL-like diagrams. Note that the reason why for the chosen benchmark point the SM-like diagrams are more important comes from the fact that $c_{t}=0.9\gg c_{g}=1/15$, compatible with current constraints. Furthermore, the higher the cut on the transverse momentum of the Higgs boson, the bigger the ratios $C/A$ and $B/A$ become, indicating again that the deviations from the SM start to become more important for high $p_{T,H}$. This stems from the fact that $d\sigma/dp_{T,H}\sim 1/p_{T,H}^{a}$ scales with $a=2$ in the full theory and with $a=1$ in the HTL Caola:2016upw . We can also see that $C_{\mathrm{HTL}}\approx C$ for all values of the cut. This is to be expected since the HTL in the SM corresponds to the HEFT with $c_{g}=2/3$, excluding those top-loops where the top quark only couples to gluons. The latter are suppressed in the HTL since they do not involve a Yukawa coupling. Thus this gives a cross-check of our computations.
@@
0 $15.0\pm 0.2$ $21.4\pm 0.5$ $30.7\pm 0.8$ $30.7\pm 0.8$
50 $9.13\pm 0.12$ $13.1\pm 0.3$ $18.9\pm 0.5$ $18.3\pm 0.5$
100 $2.86\pm 0.02$ $4.16\pm 0.07$ $6.17\pm 0.1$ $6.13\pm 0.09$
200 $(4.54\pm 0.03)\cdot 10^{-1}$ $(7.16\pm 0.12)\cdot 10^{-1}$ $(11.9\pm 0.2)\cdot 10^{-1}$ $(11.9\pm 0.1)\cdot 10^{-1}$
400 $(2.54\pm 0.01)\cdot 10^{-2}$ $(5.65\pm 0.08)\cdot 10^{-2}$ $(13.7\pm 0.1)\cdot 10^{-2}$ $(13.6\pm 0.1)\cdot 10^{-2}$
600 $(2.73\pm 0.01)\cdot 10^{-3}$ $(8.37\pm 0.11)\cdot 10^{-3}$ $(28.9\pm 0.2)\cdot 10^{-3}$ $(28.6\pm 0.1)\cdot 10^{-3}$
800 $(4.33\pm 0.02)\cdot 10^{-4}$ $(17.5\pm 0.2)\cdot 10^{-4}$ $(80.6\pm 0.5)\cdot 10^{-4}$ $(80.0\pm 0.3)\cdot 10^{-4}$
@@ 3.2 Higgs boson transverse momentum distributions
The $p_{T,H}$ distribution with a minimum $p_{T}$-cut of $30\,\mathrm{G}\mathrm{eV}$ on the jet is shown in Fig. 6, for the SM, the HTL and two HEFT benchmark points. Up to values of $p_{T,H}\approx 300~{}\mathrm{G}\mathrm{eV}$ there is no significant difference between the considered predictions, but for larger $p_{T,H}$ values the HTL shows large deviations from the SM, whereas the two HEFT parameter points lead to significantly smaller deviations. Thus it is very important to use the SM predictions with full $m_{t}$-dependence, otherwise the approximation given by the HTL could mask enhancements in the tail which are in fact BSM effects.
@@
Figure 6: Higgs boson transverse momentum distribution for two HEFT benchmark points, $(c_{t},c_{g})=(0.9,1/15)$ and $(c_{t},c_{g})=(1.1,-1/15)$, compared to the SM case and to the HTL. The bands denote 3-point scale variations around the central scale $\mu_{0}=H_{T}/2$.
@@ 4 Conclusions
We have presented results for Higgs boson production in association with one jet, combining full NLO QCD corrections with the leading operators within Higgs Effective Field Theory (HEFT), which can modify the top-Higgs Yukawa coupling and induce an effective Higgs-gluon coupling. We have taken into account constraints arising from the fact that inclusive Higgs production measurements show very good agreement with the SM prediction, assuming that subleading operators, such as the chromomagnetic operator or four-fermion operators, do not play a substantial role. We found that there are combinations of $c_{t}$ and $c_{g}$ that reproduce the inclusive Higgs production cross section at the 10% level, while changing the Higgs + jet cross section to an extent that exceeds the scale uncertainties in the highly boosted Higgs regime, i.e. for $p_{T,H}^{\mathrm{min}}\gtrsim 600~{}\mathrm{G}\mathrm{eV}$. We also found that the NLO K-factor can vary by more than 30% compared to the SM K-factor within the allowed range of $c_{t}$ and $c_{g}$. Furthermore, we showed that including the full top quark mass dependence is important to avoid that the approximation given by the HTL, leading to an enhancement in the tail of the $p_{T,H}$ distribution, is masking a BSM effect.
A more realistic study should also take into account decays of the Higgs boson as well as prospective statistical uncertainties. Furthermore, it remains to be investigated in more detail whether top mass renormalisation scheme uncertainties could swamp the effects of anomalous couplings even for highly boosted Higgs bosons. On the other hand, a better control of the top mass scheme uncertainties along the lines of Ref. Jaskiewicz:2024xkd seems feasible. In addition, the results presented in Ref. Bonciani:2022jmb for the SM case suggest that the ratio between on-shell and $\overline{\mathrm{MS}}$ renormalisation schemes for the top quark is rather flat for large $p_{T,H}$ values, while the BSM effects grow rapidly with $p_{T,H}$.
Similar considerations hold for QCD corrections beyond NLO, as the K-factors are close to constant in the large $p_{T,H}$ range. This holds for the NLO K-factors both in the HTL and in the full SM, as well as for KNNLO in the HTL Becker:2020rjp . The effects of subleading operators such as the chromomagnetic dipole operator or four-fermion operators also deserve further study, as well as electroweak corrections.
The code is available from the POWHEG-BOX-V2 website powheg-box under the process name Hjet_full_mt.
@@ A Coefficients of coupling structures
0 $15.0\pm 0.2$ $12.6\pm 0.13$ $14.2\pm 0.3$
50 $9.13\pm 0.12$ $7.53\pm 0.08$ $8.98\pm 0.21$
100 $2.86\pm 0.02$ $2.35\pm 0.02$ $2.88\pm 0.04$
200 $(4.54\pm 0.03)\cdot 10^{-1}$ $(3.69\pm 0.02)\cdot 10^{-1}$ $(4.67\pm 0.04)\cdot 10^{-1}$
400 $(2.54\pm 0.01)\cdot 10^{-2}$ $(2.04\pm 0.01)\cdot 10^{-2}$ $(2.70\pm 0.02)\cdot 10^{-2}$
600 $(2.73\pm 0.01)\cdot 10^{-3}$ $(2.19\pm 0.01)\cdot 10^{-3}$ $(2.94\pm 0.02)\cdot 10^{-3}$
800 $(4.33\pm 0.02)\cdot 10^{-4}$ $(3.46\pm 0.01)\cdot 10^{-4}$ $(4.70\pm 0.03)\cdot 10^{-4}$
@@
0 $21.4\pm 0.5$ $18.0\pm 0.3$ $22.6\pm 0.8$
50 $13.1\pm 0.3$ $10.7\pm 0.2$ $14.4\pm 0.5$
100 $4.16\pm 0.07$ $3.40\pm 0.05$ $4.62\pm 0.12$
200 $(7.16\pm 0.12)\cdot 10^{-1}$ $(5.81\pm 0.08)\cdot 10^{-1}$ $(8.05\pm 0.19)\cdot 10^{-1}$
400 $(5.65\pm 0.08)\cdot 10^{-2}$ $(4.54\pm 0.05)\cdot 10^{-2}$ $(6.46\pm 0.13)\cdot 10^{-2}$
600 $(8.37\pm 0.11)\cdot 10^{-3}$ $(6.70\pm 0.07)\cdot 10^{-3}$ $(9.62\pm 0.18)\cdot 10^{-3}$
800 $(17.5\pm 0.2)\cdot 10^{-4}$ $(14.0\pm 1.0)\cdot 10^{-4}$ $(20.1\pm 0.3)\cdot 10^{-4}$
@@
0 $30.7\pm 0.8$ $25.8\pm 0.5$ $36.3\pm 1.4$
50 $18.9\pm 0.5$ $15.4\pm 0.3$ $23.0\pm 0.8$
100 $6.17\pm 0.1$ $5.03\pm 0.08$ $7.52\pm 0.2$
200 $(11.9\pm 0.2)\cdot 10^{-1}$ $(9.74\pm 0.01)\cdot 10^{-1}$ $(14.4\pm 0.3)\cdot 10^{-1}$
400 $(13.7\pm 0.1)\cdot 10^{-2}$ $(11.2\pm 0.1)\cdot 10^{-2}$ $(16.6\pm 0.2)\cdot 10^{-2}$
600 $(28.9\pm 0.2)\cdot 10^{-3}$ $(23.4\pm 0.2)\cdot 10^{-3}$ $(34.9\pm 0.4)\cdot 10^{-3}$
800 $(80.6\pm 0.5)\cdot 10^{-4}$ $(65.2\pm 0.3)\cdot 10^{-4}$ $(97.5\pm 1.0)\cdot 10^{-4}$